if 10m of chain cost $6, calculate the cost of 1m,7m and 22m?

Answer by checkley77(12844)

6(1/10)=6*.1=$.60 for 1 m.
6(7/10)=6*.7=$4.20 for 7 m.
6(22/10)=6*2.2=$13.20 for 22 m.